OpenStudy (anonymous):

If f(x)=√8x and g(x)=x-3 whats the smallest number that is in the domain of (f o g)

OpenStudy (anonymous):

\[f\circ g(x)=f(g(x))=f)x-3)=\sqrt{x-3}\] and since you cannot take the square root of a negative number, your job is to solve \[x-3\geq 0\] for \(x\)

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

yw